Problem: Let f be a function such that f(-2)=5, f(0)=-2, and f(1)=0. Give the coordinates of the three points on the graph of y=f(x+3).
I'm not sure how to get started on this problem or where to even go once I get started. Could you write a step-by-step solution with explanations for each step? Thank you.

Answer by reviewermath(1029)

If x = -2 then y = f(-2 + 3) = f(1) = 0, the corresponding point is (-2,0).
If x = -5 then y = f(-5 + 3) = f(-2) = 5, the corresponding point is (-5,5).
If x = -3 then y = f(-3 + 3) = f(0) = -2, the corresponding point is (-3,-2)
The three points on the graph of y = f(x+3) are (-2,0), (-5,5), and (-3,-2).